OpenStudy (gabylovesyou):

What is the experimental probability of rolling a 5 on a standard die if you had 50 trials and 15 successes?

OpenStudy (gabylovesyou):

A. 35% B. 30% C. 3% D. 15%

OpenStudy (campbell_st):

the experimental probability is number of favourable outcomes over total \[P(5) = \frac{15}{50}\] convert it to a percentage

OpenStudy (gabylovesyou):

30%?

OpenStudy (unklerhaukus):

\[\frac{15}{50}=\frac{2\times15}{2\times50}=\frac{30}{100}\] yes

OpenStudy (gabylovesyou):

ok thank you!!